OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (anonymous):

x + y = 610 4x - 2y = 340 find x

OpenStudy (anonymous):

so you'll want to multiply the first equation by 2 and add them 2x + 2y = 1220 4x - 2y = 340 6x = 1560 --> x = 260